# Towards flexible perception with visual memory

### v2

@@ Abstract
Training a neural network is a monolithic endeavor, akin to carving knowledge into stone: once the process is completed, editing the knowledge in a network is nearly impossible, since all information is distributed across the network’s weights. We here explore a simple, compelling alternative by marrying the representational power of deep neural networks with the flexibility of a database. Decomposing the task of image classification into image similarity (from a pre-trained embedding) and search (via fast nearest neighbor retrieval from a knowledge database), we build a simple and flexible visual memory that has the following key capabilities: (1.) The ability to flexibly add data across scales: from individual samples all the way to entire classes and billion-scale data; (2.) The ability to remove data through unlearning and memory pruning; (3.) An interpretable decision-mechanism on which we can intervene to control its behavior. Taken together, these capabilities comprehensively demonstrate the benefits of an explicit visual memory. We hope that it might contribute to a conversation on how knowledge should be represented in deep vision models—beyond carving it in “stone” weights.
@@
1 Introduction
In the pretty diagrams on "Intro to Machine Learning” slides, an ideal ML workflow looks like this: Data collection, preprocessing, choosing a model, training, evaluation, deployment. Happy ending—the model is deployed, the users love it, and one can finally go on that well-deserved vacation and catch up on the latest AGI memes.
Until, of course, the enemy of any ideal world sets in: reality. The real world constantly keeps changing, and so do data requirements. New data and datasets become available, and existing ones become deprecated for a variety of reasons, including concerns around fairness, biases or unsafe content. Knowledge changes, and concepts drift [1; 2]: Phones and cars look different today than they did a few years ago, and different from how they will look in the future. When it comes to data, the only constant is change [3; 4; 5; 6]. Consequently, from a modeling perspective, in order to keep up with this change one would ideally want to constantly re-train or fine-tune models, which is of course not feasible. In short, as anyone who has ever deployed a model has experienced firsthand, one is constantly battling the symptoms of a single underlying cause: the fact that deep learning models have a static knowledge representation entangled in millions or billions of model parameters. We, among many others working on memory [e.g. 7; 8; 9; 10; 11; 12; 13; 14], believe that this is not a great way to represent visual knowledge for deep learning. Instead, we argue that we should build models that cleanly separate representation (how things are represented, for instance through feature embeddings) from visual memory (what is known). In short, deep learning models need a flexible visual memory: a way to explicitly utilize and edit knowledge.
In this work, we build a simple visual memory for classification and show that it has seven desirable capabilities, including the ability to flexibly add data across scales (from individual samples to classes and even billion-scale data), the ability to remove data from our model’s classification process through machine unlearning and memory pruning, and a simple, interpretable decision-mechanism on which we can intervene to control its behavior. Our main goal is to provide a compelling idea of how beneficial a flexible visual memory for deep learning can be from a variety of perspectives and capabilities. From a technical standpoint, we aim for simplicity: retrieving $k$ nearest neighbors (in an embedding feature space) along with their labels to classify a query image. This approach allows us to investigate where a simple visual memory mechanism helps, where its limitations may be, and where there might be opportunities for improvement through a more complex system. We hope that by demonstrating clear benefits from a simple visual memory, this article might contribute to a conversation on how knowledge ought to be represented in deep vision models.
Here are some highlights of this article:
  1. 1.
    Improved aggregation of retrieved memory samples: we introduce RankVoting, a power-law weighting function that surpasses previous state-of-the-art (SoftmaxVoting).
  2. 2.
    Re-ranking samples using a vision-language model achieves 88.5% top-1 ImageNet validation accuracy, improving over both DinoV2 ViT-L14 kNN and linear probing.
  3. 3.
    Flexible perception: the visual memory achieves perfect unlearning, scales to billion-scale data without additional training, and enables controlling sample influence via memory pruning.
Code and data are available here: https://github.com/google-deepmind/visual-memory
Related work. The concept of a visual memory has a long history in ML, neuroscience and psychology. In psychology, exemplar theory posits that humans recognize objects by comparing them to existing examples in visual memory [15; 16; 17; 18; 19], like the ALCOVE model [20]. In ML, prior to deep learning, instance-based learning (also known as memory-based learning) was a popular alternative to model-based learning [21; 22]. In recent years, hybrid approaches have started to combine the benefits of both approaches. For instance, deep neural network variants (model-based since they learn generalized abstractions of data) of $k$-nearest neighbor algorithms (instance-based since they compare new data to existing exemplars in memory) have been proposed with various motivations, including few-shot learning [23; 24; 25], improving adversarial robustness [26; 27; 28], medical image classification [29], confidence calibration [27], interpretability [27; 30; 31; 28], image denoising [32], retrieval-augmented learning [33; 34], anomaly and out-of-distribution detection [35; 36]. Recently, [11] tested a kNN-based visual memory up to ImageNet-scale (1.28M images), and [33; 9] applied kNN-based approaches to neural language models.
2 Building a retrieval-based visual memory for classification
Given a dataset $\mathcal{D}_{\mathsf{test}}:=\{(\tilde{\boldsymbol{x}}_{1},y_{1}),\cdots,\tilde{\boldsymbol{x}}_{n},y_{n}\}$, we want to classify each image $\tilde{\boldsymbol{x}}_{i}\in\mathcal{D}_{\mathsf{test}}$. Our classification approach consists of two steps: (i) building a visual memory, and (ii) fast nearest neighbor based inference using the visual memory.
2.1 Building a visual memory
Our visual memory retrieves (image, label) pairs from an image dataset when a query is made by directly retrieving those images that are considered similar to a test image according to a model. The model is a fixed pre-trained image encoder, meaning that no training takes place when adding information to visual memory. No copies of the dataset are stored in the visual memory. Instead, feature maps are extracted from the model based on a set of images related to the downstream classification task at hand, such as a standard training set. For our experiments, our visual memory comprises of features extracted from a dataset like the ImageNet-1K [37] training set using different encoders like DinoV2 [38] and CLIP [39]. Thus, given a pretrained image encoder, $\Phi$, and a dataset of (image, label) pairs $\mathcal{D}_{\mathsf{train}}:={(\boldsymbol{x}_{1},y_{1}),(\boldsymbol{x}_{2},y_{2}),\cdots,(\boldsymbol{x}_{N},y_{N})}$, we obtain features $\boldsymbol{z}_{i}:=\Phi(\boldsymbol{x}_{i}),\forall\boldsymbol{x}_{i}\in\mathcal{D}_{\mathsf{train}}$. Subsequently, the feature maps and corresponding label pairs are put in a database thereby creating $\mathsf{VisualMemory}:=\{(\boldsymbol{z}_{1},y_{1}),(\boldsymbol{z}_{2},y_{2}),\cdots,(\boldsymbol{z}_{N},y_{N})\}$ for classification. For both DinoV2 and CLIP, we use the last image embedding layer as a feature space.
2.2 Retrieval-based classification using visual memory
Given a query image $\tilde{\boldsymbol{x}}\in\mathcal{D}_{\mathsf{test}}$, we extract its feature map, $\tilde{\boldsymbol{z}}=\Phi(\tilde{\boldsymbol{x}})$. We then query $\mathsf{VisualMemory}$ to extract $k$ feature vectors, $\mathsf{Neighbors}({\tilde{\boldsymbol{x}}}):=\{(\boldsymbol{z}_{[1]},y_{[1]}),(\boldsymbol{z}_{[2]},y_{[2]}((,\cdots,(\boldsymbol{z}_{[k]},y_{[k]})\}$, that are closest to the query features $\tilde{\boldsymbol{z}}$ using the cosine distance, which is the default retrieval similarity measure for SSL models like DinoV2 [38]. The extracted neighbors, $\mathsf{Neighbors}({\tilde{\boldsymbol{x}}})$, are ordered by distance i.e.
@@ 2.2 Retrieval-based classification using visual memory
We then assign a weight, $w_{i}$, to each neighbour $(\boldsymbol{z}_{[i]},y_{[i]})$ and aggregate the scores for each neighbour with the same label. Finally, we assign that label to the query image with the highest aggregate score. We implemented $\mathsf{VisualMemory}$ retrieval based classification using one of the following two approaches:
@@
2. Fast and scalable nearest neighbor search: We used ScaNN [40] for accelerating nearest neighbor search at scale. Specifically, we saved the $\mathsf{VisualMemory}$ as a database and used ScaNN for fast lookup of nearest neighbors during inference. This method scales easily to billion-scale memory (cf. Section 3.3).
(a) ImageNet
(b) iNaturalist
Figure 1: Reliability of retrieved memory samples. This plot visualizes the ImageNet (left) and iNaturalist (right) top-1 validation accuracy of a single retrieved neighbor depending on the index of the neighbor (index 0: nearest neighbor). In both datasets and across models, the decrease in accuracy with increasing neighbor index follows smooth trajectories and can be approximated by a two-parameter logarithmic fit (black lines).
We mentioned earlier that we retrieve a set of neighbors, $\mathsf{Neighbors}(\tilde{\boldsymbol{x}})$ and aggregate information across them to make a classification decision. In order to understand how reliable (i.e., accurate) retrieved memory samples are from the first to the 100th neighbor, we systematically analyze neighbor reliability in Figure 1. As expected, reliability decreases as the neighbor index $k$ increases, but even at large $k$ the neighbors contain above-chance information about the ground truth class. This suggests that aggregating information across different neighbors may be beneficial to decision-making, leading to the question: What is the best aggregation strategy? We empirically study this by testing different weighting strategies for aggregation:
Plurality voting: Each neighbour in $\mathsf{Neighbors}({\tilde{\boldsymbol{x}}})$ is assigned an equal weight of 1.0. This is the classic, most simple voting method and used for instance by [11].
Distance voting: Each neighbour in $\mathsf{Neighbors}({\tilde{\boldsymbol{x}}})$ is assigned a weight based on its Cosine distance to the query image $\tilde{\boldsymbol{x}}$ i.e. $w_{i}=\exp\big{(}-\mathsf{dist}(\tilde{\boldsymbol{z}},\boldsymbol{z}_{[i]})\big{)}$. This approach has been used by [33] for nearest neighbor language models.
@@
Softmax voting: Each neighbour is assigned a weight based on the softmax function i.e. $w_{i}=\mathsf{softmax}\big{(}\mathsf{dist}(\tilde{\boldsymbol{z}},\boldsymbol{z}_{[i]}\big{)},\tau)$ where $\tau$ is the temperature. This voting method is considered state-of-the-art; for example nearest neighbor accuracies of self-supervised models are reported using this method. A temperature of $\tau=0.07$ frequently appears in literature [41; 42; 38] and is reported as a parameter “which we do not tune” in the Dino paper [42, p. 18]. We observe that performance is sensitive to this parameter; other temperatures perform worse. We therefore follow the literature in using $\tau=0.07$.
Rank voting (ours): We propose a new, simple aggregation approach wherein each neighbour is assigned a power-function weight based on its rank in the ordered set $\mathsf{Neighbors}({\tilde{\boldsymbol{x}}})$ i.e. $w_{i}=1/(\alpha+\mathsf{rank}_{i}$) where $\mathsf{rank}_{i}$ is $i$ and $\alpha$ is an offset to avoid division by zero that is set to 2.0.
(a) Existing aggregation methods, DinoV2 ViT-L14
(b) RankVoting (ours) across models
Figure 2: Aggregating information across retrieved memory samples. (left) Existing aggregation methods are overconfident in distant neighbors, resulting in the paradox of decaying ImageNet-1K accuracy with more information. The same pattern can be observed for other models and datasets in the Appendix (Figures 7 and 8). (right) This is not the case for RankVoting, our simple power-function based method which reaches higher and stable performance across models and choices of $k$.
@@
In Figure 2(a), we compare the top-1 ImageNet validation accuracy of different ranking methods as a function of number of neighbours, with the ImageNet-1K training set as the visual memory using the DinoV2/ViT-L14 model as the featurizer. Paradoxically, existing aggregation methods like plurality voting, distance-based voting, and softmax voting show decaying performance as the provided information (number of nearest neighbors) increases. This suggests that the methods are overconfident in distant neighbors, assigning them too much weight. Our simple, parameter-free rank based voting method, however, leads to an increase in performance with more neighbors until a certain $k$ after which the performance plateaus, which is the ideal scenario (Figure 2(b)). Furthermore, our rank-based voting also outperforms baselines in absolute terms; quantitative comparisons can be found in the Appendix (Tables 4, 5, 6, 7, and 8) where we also study the influence of hyperparameters (Figure 9). This indicates that a simple, power-function based method can reliably integrate information across retrieved memory samples.
@@
Gemini re-ranking. Our results above demonstrate that different aggregation strategies have a large impact on downstream performance. How far can we push the upper limit on aggregating information from different neighbors? We perform a controlled experiment using the Gemini 1.5 Flash model [43] to test this: We add the 50 nearest neighbors from DinoV2 ViT-L14 for a query image along with their labels into Gemini’s context. We then query Gemini to predict the query image’s label. This achieves 88.5% ImageNet validation accuracy, a substantial improvement over both DinoV2 ViT-L14 kNN (83.5%) and linear probing (86.3%) performance. Interestingly, the performance of the Gemini model is mainly driven by the neighbor information since it only achieves 69.6% accuracy without neighbors (when just the query image is provided to the model). The performance improvement highlights the potential of using vision-language models as a visual memory re-ranker. Given that our main goal is to explore a simple visual memory system, we mostly focus on non-Gemini ranking methods throughout our analysis.

### v3

@@ Abstract
Training a neural network is a monolithic endeavor, akin to carving knowledge into stone: once the process is completed, editing the knowledge in a network is hard, since all information is distributed across the network’s weights. We here explore a simple, compelling alternative by marrying the representational power of deep neural networks with the flexibility of a database. Decomposing the task of image classification into image similarity (from a pre-trained embedding) and search (via fast nearest neighbor retrieval from a knowledge database), we build on well-established components to construct a simple and flexible visual memory that has the following key capabilities: (1.) The ability to flexibly add data across scales: from individual samples all the way to entire classes and billion-scale data; (2.) The ability to remove data through unlearning and memory pruning; (3.) An interpretable decision-mechanism on which we can intervene to control its behavior. Taken together, these capabilities comprehensively demonstrate the benefits of an explicit visual memory. We hope that it might contribute to a conversation on how knowledge should be represented in deep vision models—beyond carving it in “stone” weights.
Machine Learning, ICML
1 Introduction
In the pretty diagrams on ”Intro to Machine Learning” slides, an ideal ML workflow looks like this: Data collection, preprocessing, choosing a model, training, evaluation, deployment. Happy ending—the model is deployed, the users love it, and one can finally go on that well-deserved vacation and catch up on the latest AGI memes.
Until, of course, the enemy of any ideal world sets in: reality. The real world constantly keeps changing, and so do data requirements. New data and datasets become available, and existing ones become deprecated for a variety of reasons, including concerns around fairness, biases or unsafe content. Knowledge changes, and concepts drift (Tsymbal, 2004; Lu et al., 2018): Phones and cars look different today than they did a few years ago, and different from how they will look in the future. When it comes to data, the only constant is change (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Zhang et al., 2023). Consequently, from a modeling perspective, in order to keep up with this change one would ideally want to constantly re-train or fine-tune models, which is not feasible. In short, as anyone who has ever deployed a model has experienced firsthand, one is constantly battling the symptoms of a single underlying cause: the fact that deep learning models have a static knowledge representation entangled in millions or billions of model parameters. We, among many others working on memory e.g. Weston et al. (2014); Chen et al. (2018); Wu et al. (2021); Iscen et al. (2022); Nakata et al. (2022); Iscen et al. (2023); Prabhu et al. (2023); Gui et al. (2024); Shao et al. (2024); Silva et al. (2024), believe that this is not a great way to represent visual knowledge for deep learning. Instead, we argue that we should build models that cleanly separate representation (how things are represented, e.g. through feature embeddings) from visual memory (what is known). In short, deep learning models need a flexible visual memory: a way to explicitly utilize and edit knowledge.
In this work, we build a simple visual memory for classification and show that it has seven desirable capabilities, including the ability to flexibly add data across scales (from individual samples to classes and even billion-scale data), the ability to remove data from our model’s classification process through machine unlearning and memory pruning, and a simple, interpretable decision-mechanism on which we can intervene to control its behavior. Our main goal is to provide a compelling idea of how beneficial a flexible visual memory for deep learning can be from a variety of perspectives and capabilities. From a technical standpoint, we aim for simplicity: retrieving $k$ nearest neighbors (in an embedding feature space) along with their labels to classify a query image. This approach allows us to investigate where a simple visual memory mechanism helps, where its limitations may be, and where there might be opportunities for improvement through a more complex system. We hope that by demonstrating clear benefits from a simple visual memory, this article might contribute to a conversation on how knowledge ought to be represented in deep vision models. Our contributions are as follows:
  1. 1.
    Our main contribution is to demonstrate that a visual memory based system enables flexible capabilities, making progress towards some of deep learning’s grand challenges: In specific settings, unlearning with perfect guarantees which is currently considered a major open problem in deep learning, an improved understanding of how decisions relate to datapoints through an interpretable decision-mechanism enabling data attribution, and controlling sample influence through memory pruning. These capabilities highlight the promise of separating knowledge from representation through a visual memory.
  2. 2.
    A simple visual memory performs well at scale: On the methodological side, we aim for simplicity. Instead of re-inventing the wheel, we build on established building blocks from the literature like SSL features, kNN classification and fast, scalable similarity search. We contribute technical improvements like RankVoting and VLM re-ranking, achieving 88.5% top-1 ImageNet validation accuracy which improves over both DinoV2 ViT-L14 kNN and linear probing.
Figure 1: Reliability of retrieved memory samples. This plot visualizes the ImageNet top-1 validation accuracy of a single retrieved neighbor depending on the index of the neighbor (index 0: nearest neighbor). The decrease in accuracy with increasing neighbor index follows smooth trajectories and can be approximated by a two-parameter logarithmic fit (black lines). More detailed plot in Appendix H.
We argue that the way current deep learning models represent knowledge (static knowledge representation, hard to update, hard to unlearn, hard to understand how a decision is made) is problematic. As an alternative, we built a working proof-of-concept: By building on the long history of nearest neighbor methods, and “marrying” them with a powerful deep learning representation (such as SSL features from DinoV2) and a billion-scale visual memory.
Related work. The concept of a visual memory has a long history in ML, neuroscience and psychology. In psychology, exemplar theory posits that humans recognize objects by comparing them to existing examples in visual memory (Medin & Schaffer, 1978; Nosofsky, 1986; Dopkins & Gleason, 1997; Jäkel et al., 2008; Nosofsky, 2011), like the ALCOVE model (Kruschke, 2020). In ML, prior to deep learning, instance-based learning (also known as memory-based learning) was a popular alternative to model-based learning (Aha et al., 1991; Quinlan, 1993). For instance, Turk & Pentland (1991) used nearest neighbor methods to classify faces, and Sivic & Zisserman (2003) build a visual memory inspired by text retrieval for object retrieval from videos. In recent years, hybrid approaches have started to combine the benefits of both approaches. Deep neural network variants (model-based since they learn generalized abstractions of data) of $k$-nearest neighbor algorithms (instance-based since they compare new data to existing exemplars in memory) have been proposed with various motivations, including few-shot learning (Wang et al., 2019b; Yang et al., 2020; Bari et al., 2021), improving adversarial robustness (Sitawarin & Wagner, 2019; Papernot & McDaniel, 2018; Rajani et al., 2020), medical image classification (Zhuang et al., 2020), confidence calibration (Papernot & McDaniel, 2018), interpretability (Papernot & McDaniel, 2018; Wallace et al., 2018; Lee et al., 2020; Rajani et al., 2020), image denoising (Plötz & Roth, 2018), retrieval-augmented learning (Khandelwal et al., 2019; Drozdov et al., 2022), anomaly and out-of-distribution detection (Bergman et al., 2020; Sun et al., 2022). Recently, Nakata et al. (2022) tested a kNN-based visual memory up to ImageNet-scale (1.28M images), and Khandelwal et al. (2019); Wu et al. (2021) applied kNN-based approaches to neural language models. In contrast, we scale visual memory to the billion scale, improve ranking, and show systematic benefits across different tasks.
2 Building a retrieval-based visual memory for classification
Given a dataset $\mathcal{D}_{\mathsf{test}}:=\{(\tilde{\boldsymbol{x}}_{1},y_{1}),\cdots,\tilde{\boldsymbol{x}}_{n},y_{n}\}$, we want to classify each image $\tilde{\boldsymbol{x}}_{i}\in\mathcal{D}_{\mathsf{test}}$. Our classification approach consists of two steps: (i) building a visual memory, and (ii) using it for fast nearest neighbor based inference.
(a) Existing aggregation methods, DinoV2 ViT-L14
(b) RankVoting across models
Figure 2: Aggregating information across retrieved memory samples. (left) Existing aggregation methods are overconfident in distant neighbors, resulting in the paradox of decaying ImageNet-1K accuracy with more information. The same pattern is also seen for other models and datasets in the Appendix (Figures 7 and 8). (right) This is not the case for RankVoting, a power-function based method which reaches higher and stable performance across models and choices of $k$.
2.1 Building a visual memory
Our visual memory retrieves (image, label) pairs from an image dataset when a query is made by directly retrieving those images that are considered similar to a test image according to a model. The model is a fixed pre-trained image encoder, meaning that no training takes place when adding information to visual memory. No copies of the dataset are stored in the visual memory. Instead, feature maps are extracted from the model based on a set of images related to the downstream classification task at hand, such as a standard training set. For our experiments, our visual memory comprises of features extracted from a dataset like the ImageNet-1K (Russakovsky et al., 2015) training set or JFT (Zhai et al., 2022) using different encoders like DinoV2 (Oquab et al., 2023) and CLIP (Radford et al., 2021). Thus, given a pretrained image encoder, $\Phi$, and a dataset of (image, label) pairs $\mathcal{D}_{\mathsf{train}}:={(\boldsymbol{x}_{1},y_{1}),(\boldsymbol{x}_{2},y_{2}),\cdots,(\boldsymbol{x}_{N},y_{N})}$, we obtain features $\boldsymbol{z}_{i}:=\Phi(\boldsymbol{x}_{i}),\forall\boldsymbol{x}_{i}\in\mathcal{D}_{\mathsf{train}}$. Subsequently, the feature maps and corresponding label pairs are put in a database thereby creating $\mathsf{VisualMemory}:=\{(\boldsymbol{z}_{1},y_{1}),(\boldsymbol{z}_{2},y_{2}),\cdots,(\boldsymbol{z}_{N},y_{N})\}$ for classification. For both DinoV2 and CLIP, we use the last image embedding layer as a feature space.
2.2 Retrieval-based classification using visual memory
@@
We then assign a weight, $w_{i}$, to each neighbour $(\boldsymbol{z}_{[i]},y_{[i]})$ and aggregate the scores for each neighbour with the same label. Finally, we assign that label to the query image with the highest aggregate score. We implemented retrieval based classification using one of the following two approaches:
@@
2. Fast and scalable nearest neighbor search: We used ScaNN (Guo et al., 2020) for accelerating nearest neighbor search at scale. Specifically, we saved the $\mathsf{VisualMemory}$ as a database and used ScaNN for fast lookup of nearest neighbors during inference. Storing features requires only about 1–3% of the space of storing the dataset itself and search latency is 500–600 QPS at perfect recall for 1M features, thus approximately 2 milliseconds per query due to parallelization which can be done on CPUs. See Appendix K for details on latency and storage. This method scales easily to billion-scale memory (cf. Section 3.3).
[…]
@@
Softmax voting: Each neighbour is assigned a weight based on the softmax function i.e. $w_{i}=\mathsf{softmax}\big{(}\mathsf{dist}(\tilde{\boldsymbol{z}},\boldsymbol{z}_{[i]}\big{)},\tau)$ where $\tau$ is the temperature. This voting method is considered state-of-the-art; for example nearest neighbor accuracies of self-supervised models are reported using this method. A temperature of $\tau=0.07$ frequently appears in literature (Wu et al., 2018; Caron et al., 2021; Oquab et al., 2023) and is reported as a parameter “which we do not tune” in the Dino paper (Caron et al., 2021, p. 18). We observe that performance is sensitive to this parameter; other temperatures perform worse. We therefore follow the literature in using $\tau=0.07$.
[…]
@@
In Figure 2(a), we compare the top-1 ImageNet validation accuracy of different ranking methods as a function of number of neighbours, with the ImageNet-1K training set as the visual memory using the DinoV2/ViT-L14 model as the featurizer. Paradoxically, existing aggregation methods like plurality voting, distance-based voting, and softmax voting show decaying performance as the provided information (number of nearest neighbors) increases. This suggests that the methods are overconfident in distant neighbors, assigning them too much weight. Our simple, parameter-free rank based voting method, however, leads to an increase in performance with more neighbors until a certain $k$ after which the performance plateaus, which is the ideal scenario (Figure 2(b)). Furthermore, rank-based voting also outperforms baselines in absolute terms; quantitative comparisons can be found in the Appendix (Tables 4, 5, 6, 7 and 8) where we also study the influence of hyperparameters (Figure 9). This indicates that a simple, power-function based method can reliably integrate information across retrieved memory samples.
@@
Gemini re-ranking. Our results above demonstrate that different aggregation strategies have a large impact on downstream performance. How far can we push the upper limit on aggregating information from different neighbors? We perform a controlled experiment using the Gemini 1.5 Flash model (Reid et al., 2024) to test this: We add the 50 nearest neighbors from DinoV2 ViT-L14 for a query image along with their labels into Gemini’s context. We then query Gemini to predict the query image’s label. This achieves 88.5% ImageNet validation accuracy, a substantial improvement over both DinoV2 ViT-L14 kNN (83.5%) and linear probing (86.3%) performance. Interestingly, Gemini’s performance is mainly driven by the neighbor information through in-context learning since it only achieves 69.6% accuracy without neighbors (when just the query image is provided to the model). The performance improvement highlights the potential of using vision-language models as a visual memory re-ranker. Given that our main goal is to explore a simple visual memory system, we mostly focus on non-Gemini ranking methods throughout our analysis.